# A unified linear intrinsic alignment model for elliptical and disc galaxies and the resulting ellipticity spectra

### v1

@@ 2.4.2 Tidal effect on moments of the brightness distribution
which directly reflects magnitude and orientation of the tidal shear field $\partial^{2}_{ij}\Phi$, with the susceptibility $S_{abij}$ as the constant of proportionality: It includes the virial velocity $\sigma_{\mathrm{vir}}^{2}$, and the details of the brightness distribution $\rho_{0}$. For an elliptical, $S_{abij}$ is fundamentally an area of size $R^{2}$ for a Sérsic profile with scale $R$, and a numerical factor resulting from the integration that depends strongly on Sérsic-index $n$ (Ghosh et al.(2021)Ghosh, Durrer, and Schaefer; Giesel et al.(2022a)Giesel, Ghosh, and Schaefer; Giesel et al.(2022b)Giesel, Ghosh, and Schäfer): Typical numerical prefactors to $R^{2}$ resulting from the integration of the Sérsic-profile are $S\simeq 10$ for exponential profiles with $n=1$ up to $S\simeq 10^{3}$ for de Vaucouleurs-profiles with $n=4$. The reaction of a galactic disc to a tidal shear field can in principle be encapsulated by the same relation, with $\upsilon_{\mathrm{circ}}^{2}$ replacing $\sigma_{\mathrm{vir}}^{2}$, and absorbing the averaging in eqn. (24) into $S_{abij}$.
@@ 2.4.3 Consistency relations
For spiral galaxies we expect an identical argument to hold, with the choice of an exponential profile typical for galactic discs, and a radius-averaged ellipticity, as deduced in the previous chapter. But there are two major differences: Tidal alignments of orbits in the galactic disc are strongly radius-dependent, necessitating the averaging process outlined in eqn. (24). Secondly, the squared Sérsic-radius, which is a sensible choice for the alignment parameter of elliptical galaxies as it reproduces the alignment amplitude and corresponds to the physical picture of a tidally distorted virialised object, is remarkably similar for spiral and elliptical galaxies of similar mass. Therefore, it cannot be the alignment parameter for a disc. If one stipulates that $SR^{2}=r_{s}^{2}$ for disc galaxies and $SR^{2}=r_{\mathrm{vir}}^{2}$ for ellipticial galaxies, the missing link between the dynamical properties of the dark matter halo (encapsulated in $r_{s}$ and $r_{\mathrm{vir}}$) and the light distribution of the luminous component (described by $R$ and $n$) can potentially be established: Numerically, this would imply for a Milky Way-sized halo of $\sim 10^{12}M_{\odot}/h$ with a virial radius of $\sim 100~{}\mathrm{kpc}/h$ (Dehnen et al.(2006)Dehnen, McLaughlin, and Sachania), an NFW-scale radius $r_{s}\sim 10~{}\mathrm{kpc}/h$ (Lin and Li(2019)) and a typical Sérsic-radius of a few kpc (Goodwin et al.(1997)Goodwin, Gribbin, and Hendry; Goodwin et al.(1998)Goodwin, Gribbin, and Hendry), which would not be too dissimilar for an elliptical of the same mass.
@@ 2.4.4 Observational consequences
Figure 3: Scaling of the concentration parameter $c(M_{\mathrm{vir}},z)$ with virial halo mass $M_{\mathrm{vir}}$ and redshift $z$ for two common halo profile families (top panel) and dependence of the squared velocity ratio $\sigma^{2}/\upsilon_{\mathrm{circ}}^{2}$ (bottom panel), both derived with the fitting formula from 2018.
@@ 2.4.5 Tidal fields in the large-scale structure
and the tidal field exerted by the large-scale structure is again weak, $\omega^{2}/\Omega^{2}=\Delta/3\gg 1$. On both cases, the galaxies are sufficiently strongly bound such that typical tidal fields are weak perturbations to their own gravitational fields. We suspect that the inverse density scale introduced in the common parameterisation of the linear alignment model for elliptical galaxies (Hirata and Seljak(2004); Brown et al.(2002)Brown, Taylor, Hambly, and Dye; Blazek et al.(2019)Blazek, MacCrann, Troxel, and Fang; Blazek et al.(2011)Blazek, McQuinn, and Seljak; Blazek et al.(2015)Blazek, Vlah, and Seljak)
@@
with $\tilde{C}_{1}$ fixed to $5\times 10^{-14}(\mathrm{Mpc}/h)^{3}/(M_{\odot}/h)$ effectively reproduces with the background density $\rho_{\mathrm{crit}}\Omega_{m}$ the virial overdensity $\Delta$, as it determines the ratio between internal and external gravitational fields.
@@ 2.5 Tidal fields in 3 dimensions
where a suitable basis $\lambda^{(n)}_{ij}$ could be provided by the symmetric subset of Gell-Mann matrices (as a generalisation of the Pauli-matrices used in the analogous expansion in eqns. 6 and 7), which form together with the unit matrix a basis of all tidal shear fields $\partial_{i}\partial_{j}\Phi$. There are six symmetric Gell-Mann matrices, and the effect of two of them, specifically $\lambda^{1}$ and $\lambda^{(3)}$, was already investigated above, as they contain the Pauli-matrices in their top left $2\times 2$ submatrix:
@@
The unit matrix, which is likewise an element of the Gell-Mann matrices, would cause an isotropic change in size of the galactic disc (c.f. Appendix Appendix: Isotropic increase in size). The two Gell-Mann matrices $\lambda^{(4)}$ and $\lambda^{(6)}$,
@@ 3 Ellipticity spectra
The primary observable of weak lensing surveys are (tomographic) ellipticity spectra, which map out line of-sight-integrated fluctuations of the tidal gravitational field in a wide redshift range reaching out beyond unity. As a subdominant contribution to the lensing-induced ellipticity spectra there are intrinsic alignments of elliptical galaxies, which reflect redshift-distribution averaged correlations of tidal shear fields in the respective redshift bin. With the results from the previous chapters, namely that disc galaxies follow the same linear alignment model as ellipticals albeit with a smaller alignment parameter, one can relate their intrinsic ellipticity spectra by an overall scaling. Numerically, we follow the analysis outlined by Ghosh et al.(2021)Ghosh, Durrer, and Schaefer, to which we refer the reader.
@@
Angular ellipticity spectra for lensing $C_{ij}^{\gamma\gamma}(\ell)$, for intrinsic alignments $C_{ij}^{\epsilon\epsilon}(\ell)$ and for their cross-correlation $C_{ij}^{\gamma\epsilon}(\ell)$ are shown in Fig. 5, for a Euclid-like survey with median redshift of 0.9. For clearer visualisation, we choose to show a simplified setup with only three bins, with equal fractions of the total galaxy population. While there are no cross-correlations across different redshifts bins in intrinsic alignments, cross-correlations of lensing do exist, but are excluded in the plots in order not to overcrowd the figure. Parallel to the $GI$ and $II$-spectra for elliptical galaxies we show the corresponding spectra of disc galaxies as well, both weighted by the relative fraction of elliptical galaxies (making up roughly a third of the galaxy population) and disc galaxies (accounting for the remaining two thirds). The shape noise term assumes a standard value of $\sigma_{\epsilon}=0.4$ for an individual galaxy, scaled down with $n=4.727\times 10^{8}$ galaxies per steradian.
@@
Figure 5: Angular ellipticity spectra for a simplified three-bin tomographic survey with Euclid’s redshift distribution, comparing weak gravitational lensing with intrinsic alignments of elliptical galaxies, and in particular with alignments of disc galaxies according to a quadratic and to a linear alignment model. We assume that spirals make up two thirds of the galaxy population, and ellipticals the remaining third.
@@
The ratios $C^{\epsilon\epsilon}_{ij}(\ell)/C^{\gamma\gamma}_{ij}(\ell)$ and $C^{\gamma\epsilon}_{ij}(\ell)/C^{\gamma\gamma}_{ij}(\ell)$ between the intrinsic alignment spectra and the weak lensing spectrum are shown in Fig. 6, again for only three bins and for the choice $i=j$. Most importantly, one can see by how much an angular-momentum based, quadratic alignment model would overpredict the effect (Schäfer and Merkel(2017)), and how the linear alignment model for spiral galaxies evades detection in earlier data sets (Heymans et al.(2013)). But most importantly, the dynamical model suggests that the alignment amplitudes of elliptical and spiral galaxies should be related, as the alignment parameter reflects effectively the area of the luminous component of the systems in both cases, suggesting the question whether marked correlation functions which combine intrinsic ellipticity correlations with galaxy properties such as Sérsic-indices or Sérsic-radii can isolate more information about the alignment mechanisms.
@@
Figure 6: Ratios between the angular ellipticity spectra of intrinsic alignments and weak gravitational lensing analogous to Fig. 5.
@@
Starting with the numbers derived by Ghosh et al.(2021)Ghosh, Durrer, and Schaefer, who derive estimates on the observability of ellipticity-lensing correlation in aligned ellipticals in a Euclid-like survey: For ellipticals making up $q=1/3$ of the galaxy population and an alignment parameter $D_{IA}=-(10^{-5}~{}\mathrm{Mpc}/h)^{2}$, a Euclid-like redshift distribution with a median redshift of $0.9$, 5 tomographic bins and integrating up all multipoles from $\ell=10$ to $\ell=3000$ yields a statistical significance $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 10\ldots 15$ including effects on nonlinear structure formation (and correspondingly higher tidal shear fields) on small scales. Assuming that the measurement is limited by shape noise and by cosmic variance from weak gravitational lensing, one can use the proportionality of the spectra to scale down the significance for ellipticity correlations of disc galaxies: Relative to the $GI$-spectrum for ellipticals which is measurable at $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 15$, the $GI$-spectrum for spirals should generate a signal at a significance lower by $\sqrt{(1-q)/q/\alpha}\simeq 0.7$. We assume the ratio of alignment parameters $\alpha$ a value of $4$, intermediate between Hilbert et al.(2017)Hilbert, Xu, Schneider, Springel, Vogelsberger, and Hernquist and Jagvaral et al.(2022)Jagvaral, Singh, and Mandelbaum. Relative to the $II$-spectrum of elliptical galaxies can be measured at a signficiance of about $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 10$, the $II$-spectrum of spirals alone should generate a significance lower by $(1-q)/q/\alpha\simeq 0.5$, and the $II$-type cross-correlation between spirals and ellipticals would be at significances lower by $\sqrt{(1-q)/\alpha}\simeq 0.4$. This effectively puts all spectra within reach of a Euclid-type survey.
@@ 4 Summary and discussion
    With the alignment of both elliptical and spiral galaxies following a linear alignment model and both being determined by the external tidal gravitational field, the question of how the relative alignments are determined by galaxy properties arises naturally. Elliptical galaxies react in a homologous way to tidal fields, whereas the tidally-induced change in shape of spirals depends on radius. The linear alignment model would now stipulate that the alignment of spirals is weaker as the discs are smaller than the virial radius of the host halo. At the NFW-scale radius, the ellipticity of a galactic disc is $c(M_{\mathrm{vir}},z)^{2}$ times smaller than that of an elliptical, with $c(M_{\mathrm{vir}},z)$ being the NFW-concentration parameter. Averaging the ellipticity over all radii, weighted with an exponential brightness distribution alleviates this number and sets the alignments strength of spirals to be about an order of magnitude smaller than that of ellipticals.
@@
    The physical picture behind linear alignments of disc galaxies suggests a number of observational relations that could be tested with data: The alignment should be stronger with higher Sérsic-index, and higher for systems with larger discs. There should be a scaling of the alignment with the mass of the system, driven by the change in NFW-concentration parameter $c(M_{\mathrm{vir}},z)$, too: As low-mass systems tend to be more centrally concentrated, their alignment should be weaker. Whether empirically introduced laws to allow for redshift evolution can be replaced by the known evolution of the NFW-concentration parameter $c(M_{\mathrm{vir}},z)$: Here, the alignment would be highest for high $z$ and decrease towards the current epoch. To what extent evolution of alignments with redshift and their scaling with mass are driven by the dependence of the NFW-concentration parameter and what additional astrophysical effects could play a role would be an interesting question for surveys and for simulations.
@@
    A further interesting consequence of the linear alignment model are changes in shape: Both ellipticals and spirals should tend to be larger in low-density regions. Flexion-like distortions from higher-order tidal shear fields on galactic discs should exist but would be very small, given the result that the effect induced by $\partial^{3}\Phi$ on ellipticals as investigated in Giesel et al.(2022a)Giesel, Ghosh, and Schaefer is already small. But physically, these distortions would be able to warp a galactic disc, which would constitute a link between alignments and disc dynamics.
@@
    Ellipticity spectra resulting from the linear alignment model naturally resemble those of elliptical galaxies as they are related just by a constant prefactor. With their smaller amplitudes ellipticity correlations of spiral galaxies evade constraints on alignments derived from current weak lensing data and do not produce signals that can be measured at reasonable significances. More important, though, are the much smaller amplitudes of the linear alignment model in spiral galaxies compared to angular-momentum based models: Typically, the spectra are smaller by more than an order of magnitude at high multipoles. Additionally, the phenomenology of the spectra is quite different: In contrast to the quadratic, angular-momentum based alignment model which predicts no cross-correlation between the shapes of elliptical and spiral galaxies, the linear alignment model would predict cross-correlations between the shapes of the two galaxy types in a very natural way. Moreover, the linear alignment model would produce pure $E$-mode correlations in the ellipticity field, which would apply likewise for spiral and elliptical galaxies. Again, the quadratic model would be markedly different, with a high amplitude $B$-mode correlation being predicted. Lastly, one would expect a nonzero cross-correlation between intrinsic shapes of both ellipticals and spirals with gravitational lensing shear, with the typical anti-correlations present in both galaxy types. Linking tidal alignment to galaxy properties one should expect a dependence on Sérsic-index as more centrally concentrated objects would show weaker alignments. Estimates suggest that all intrinsic ellipticity spectra are within reach of Euclid.
Discs galaxies regularly have galactic bulges at their centres, which show analogous alignments in simulations by Jagvaral et al.(2022)Jagvaral, Singh, and Mandelbaum, up to the fact that they tend to have rather high Sérsic indices close to $n=4$, are rather strongly bound with high velocity dispersions $\sigma$ and are rather compact with a small Sérsic radii $R$; numbers of $\sigma=140~{}\mathrm{km}/\mathrm{s}$ and $R=1/2~{}\mathrm{kpc}$ have been reported for Milky Way (Valenti et al.(2018)Valenti, Zoccali, Mucciarelli, Gonzalez, Surot, Minniti, Rejkuba, Pasquini, Fiorentino, Bono et al.). Their amplitude seems to be similar to that of elliptical galaxies, and is much stronger than that of the host disc galaxy. Whether this is a consequence of the homologous distortion of virial systems remains to be investigated. We intend to develop the topic of tidally distorted galactic discs and the resulting ellipticity correlations further by considering the full 3-dimensional tidal field, which would effectively bridge between the concept of disc warps and intrinsic alignments, or perhaps even link alignments with tidally-induced bar formation in spiral galaxies.

### v2

@@ 2.4.2 Tidal effect on moments of the brightness distribution
which directly reflects magnitude and orientation of the tidal shear field $\partial^{2}_{ij}\Phi$, with the susceptibility $S_{abij}$ as the constant of proportionality: It includes the virial velocity $\sigma_{\mathrm{vir}}^{2}$, and the details of the brightness distribution $\rho_{0}$. For an elliptical, $S_{abij}$ is fundamentally an area of size $R^{2}$ for a Sérsic profile with scale $R$, and a numerical factor resulting from the integration that depends strongly on Sérsic-index $n$ (Ghosh et al., 2021; Giesel et al., 2022a, b): Typical numerical prefactors to $R^{2}$ resulting from the integration of the Sérsic-profile are $S\simeq 10$ for exponential profiles with $n=1$ up to $S\simeq 10^{3}$ for de Vaucouleurs-profiles with $n=4$. The reaction of a galactic disc to a tidal shear field can in principle be encapsulated by the same relation, with $\upsilon_{\mathrm{circ}}^{2}$ replacing $\sigma_{\mathrm{vir}}^{2}$, and absorbing the averaging in eqn. (24) into $S_{abij}$.
@@ 2.4.3 Consistency relations
For spiral galaxies we expect an identical argument to hold, with the choice of an exponential profile typical for galactic discs, and a radius-averaged ellipticity, as deduced in the previous chapter. But there are two major differences: Tidal alignments of orbits in the galactic disc are strongly radius-dependent, necessitating the averaging process outlined in eqn. (24). Secondly, the squared Sérsic-radius, which is a sensible choice for the alignment parameter of elliptical galaxies as it reproduces the alignment amplitude and corresponds to the physical picture of a tidally distorted virialised object, is remarkably similar for spiral and elliptical galaxies of similar mass. Therefore, it cannot be the alignment parameter for a disc. If one stipulates that $SR^{2}=r_{s}^{2}$ for disc galaxies and $SR^{2}=r_{\mathrm{vir}}^{2}$ for elliptical galaxies, the missing link between the dynamical properties of the dark matter halo (encapsulated in $r_{s}$ and $r_{\mathrm{vir}}$) and the light distribution of the luminous component (described by $R$ and $n$) can potentially be established: Numerically, this would imply for a Milky Way-sized halo of $\sim 10^{12}M_{\odot}/h$ with a virial radius of $\sim 100~{}\mathrm{kpc}/h$ (Dehnen et al., 2006), an NFW-scale radius $r_{s}\sim 10~{}\mathrm{kpc}/h$ (Lin and Li, 2019) and a typical Sérsic-radius of a few kpc (Goodwin et al., 1997, 1998), which would not be too dissimilar for an elliptical of the same mass.
@@ 2.4.5 Tidal fields in the large-scale structure
and the tidal field exerted by the large-scale structure is again weak, $\omega^{2}/\Omega^{2}=\Delta/3\gg 1$. On both cases, the galaxies are sufficiently strongly bound such that typical tidal fields are weak perturbations to their own gravitational fields. We suspect that the inverse density scale introduced in the common parameterisation of the linear alignment model for elliptical galaxies (Hirata and Seljak, 2004; Brown et al., 2002; Blazek et al., 2019, 2011, 2015)
@@
with $\tilde{C}_{1}$ fixed to $5\times 10^{-14}(\mathrm{Mpc}/h)^{3}/(M_{\odot}/h)$ effectively reproduces with the background density $\rho_{\mathrm{crit}}\Omega_{m}$ the virial overdensity $\Delta$, as it determines the ratio between internal and external gravitational fields. This has indeed been proven to be true and mathematically derived previously in (Hirata and Seljak, 2004).
@@ 2.5 Tidal fields in 3 dimensions
where a suitable basis $\lambda^{(n)}_{ij}$ could be provided by the symmetric subset of Gell-Mann matrices (as a generalisation of the Pauli-matrices used in the analogous expansion in eqns. 6 and 7), which form together with the unit matrix a basis of all tidal shear fields $\partial_{i}\partial_{j}\Phi$. There are six symmetric Gell-Mann matrices, and the effect of two of them, specifically $\lambda^{(1)}$ and $\lambda^{(3)}$, was already investigated above, as they contain the Pauli-matrices in their top left $2\times 2$ submatrix:
@@
The unit matrix, which is likewise an element of the Gell-Mann matrices, would cause an isotropic change in size of the galactic disc (c.f. Appendix). The two Gell-Mann matrices $\lambda^{(4)}$ and $\lambda^{(6)}$,
@@ 3 Ellipticity spectra
The primary observable of weak lensing surveys are (tomographic) ellipticity spectra, which map out line of-sight-integrated fluctuations of the tidal gravitational field in a wide redshift range reaching out beyond unity. As a subdominant contribution to the lensing-induced ellipticity spectra there are intrinsic alignments of elliptical galaxies, which reflect redshift-distribution averaged correlations of tidal shear fields in the respective redshift bin. With the results from the previous chapters, namely that disc galaxies follow the same linear alignment model as ellipticals albeit with a smaller alignment parameter, one can relate their intrinsic ellipticity spectra by an overall scaling. Numerically, we follow the analysis outlined by Ghosh et al. (2021), to which we refer the reader.
@@
Angular ellipticity spectra for lensing $C_{ij}^{\gamma\gamma}(\ell)$, for intrinsic alignments $C_{ij}^{\epsilon\epsilon}(\ell)$ and for their cross-correlation $C_{ij}^{\gamma\epsilon}(\ell)$ are shown in Fig. 5, for an Euclid-like survey with median redshift of 0.9. For clearer visualisation, we choose to show a simplified setup with only three bins, with equal fractions of the total galaxy population. While there are no cross-correlations across different redshifts bins in intrinsic alignments, cross-correlations of lensing do exist, but are excluded in the plots in order not to overcrowd the figure. Parallel to the $GI$ and $II$-spectra for elliptical galaxies we show the corresponding spectra of spiral galaxies as well, both weighted by the relative fraction of elliptical galaxies (making up roughly a third of the galaxy population) and spiral galaxies (accounting for the remaining two thirds). The shape noise term assumes a standard value of $\sigma_{\epsilon}=0.4$ for an individual galaxy, scaled down with $n=4.727\times 10^{8}$ galaxies per steradian.
@@
Figure 5: Angular ellipticity spectra for a simplified three-bin tomographic survey with Euclid’s redshift distribution, comparing weak gravitational lensing with intrinsic alignments of elliptical galaxies, and in particular with alignments of spiral galaxies according to a quadratic and to a linear alignment model. We assume that spirals make up two thirds of the galaxy population, and ellipticals the remaining third.
@@
The ratios $C^{\epsilon\epsilon}_{ij}(\ell)/C^{\gamma\gamma}_{ij}(\ell)$ and $C^{\gamma\epsilon}_{ij}(\ell)/C^{\gamma\gamma}_{ij}(\ell)$ between the intrinsic alignment spectra and the weak lensing spectrum are shown in Fig. 6, again for only three bins and for the choice $i=j$. Most importantly, one can see by how much an angular-momentum based, quadratic alignment model would overpredict the effect (Schäfer and Merkel, 2017), and how the linear alignment model for spiral galaxies evades detection in earlier data sets (Heymans et al., 2013). But most importantly, the dynamical model suggests that the alignment amplitudes of elliptical and spiral galaxies should be related, as the alignment parameter reflects effectively the area of the luminous component of the systems in both cases. This suggests the question whether marked correlation functions which combine intrinsic ellipticity correlations with galaxy properties such as Sérsic-indices or Sérsic-radii can isolate more information about the alignment mechanisms.
@@
Figure 6: Ratios between the angular ellipticity spectra of intrinsic alignments and weak gravitational lensing analogous to Fig. 5, for three bins and for $i=j$. It is clear that there is a large overprediction in case of the quadratic alignment model for spiral galaxies.
@@
We start with the numbers derived by Ghosh et al. (2021), who derive estimates on the observability of ellipticity-lensing correlation in aligned ellipticals in an Euclid-like survey. For ellipticals making up $q=1/3$ of the galaxy population and an alignment parameter $D_{IA}=-(10^{-5}~{}\mathrm{Mpc}/h)^{2}$, an Euclid-like redshift distribution with a median redshift of $0.9$, 5 tomographic bins and integrating up all multipoles from $\ell=10$ to $\ell=3000$ yields a statistical significance $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 10\ldots 15$, where $\Sigma$ is the signal-to-noise ratio and $f_{\mathrm{sky}}$ is the sky coverage of the survey, including effects on nonlinear structure formation (and correspondingly higher tidal shear fields) on small scales. Assuming that the measurement is limited by shape noise and by cosmic variance from weak gravitational lensing, one can use the proportionality of the spectra to scale down the significance for ellipticity correlations of disc galaxies: Relative to the $GI$-spectrum for ellipticals which is measurable at $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 15$, the $GI$-spectrum for spirals should generate a signal at a significance lower by $\sqrt{(1-q)/q/\alpha}\simeq 0.7$. We assume the ratio of alignment parameters $\alpha$ a value of $4$, intermediate between Hilbert et al. (2017) and Jagvaral et al. (2022). Relative to the $II$-spectrum of elliptical galaxies can be measured at a signficiance of about $\Sigma/\sqrt{f_{\mathrm{sky}}}\simeq 10$, the $II$-spectrum of spirals alone should generate a significance lower by $(1-q)/q/\alpha\simeq 0.5$, and the $II$-type cross-correlation between spirals and ellipticals would be at significances lower by $\sqrt{(1-q)/\alpha}\simeq 0.4$. This effectively puts all spectra within reach of an Euclid-type survey.
@@ 4 Summary and discussion
    With the alignment of both elliptical and spiral galaxies following a linear alignment model and both being determined by the external tidal gravitational field, the question of how the relative alignments are determined by galaxy properties arises naturally. Elliptical galaxies react in a homologous way to tidal fields, whereas the tidally-induced change in shape of spirals depends on radius. The linear alignment model would now stipulate that the alignment of spirals is weaker as the discs are smaller than the virial radius of the host halo. At the NFW-scale radius, the ellipticity of a galactic disc is $c(M_{\mathrm{vir}},z)^{2}$ times smaller than that of an elliptical, with $c(M_{\mathrm{vir}},z)$ being the NFW-concentration parameter. Averaging the ellipticity over all radii, weighted with an exponential brightness distribution, alleviates this number and sets the alignments strength of spirals to be about an order of magnitude smaller than that of ellipticals.
@@
    The physical picture behind linear alignments of disc galaxies suggests a number of observational relations that could be tested with data: The alignment should be stronger with higher Sérsic-index, and higher for systems with larger discs. There should be a scaling of the alignment with the mass of the system, driven by the change in NFW-concentration parameter $c(M_{\mathrm{vir}},z)$, too: As low-mass systems tend to be more centrally concentrated, their alignment should be weaker. It can also be checked whether empirically introduced laws to allow for redshift evolution can be replaced by the known evolution of the NFW-concentration parameter $c(M_{\mathrm{vir}},z)$: Here, the alignment would be highest for high $z$ and decrease towards the current epoch. To what extent evolution of alignments with redshift and their scaling with mass are driven by the dependence of the NFW-concentration parameter and what additional astrophysical effects could play a role, would be an interesting question for surveys and for simulations.
@@
    A further interesting consequence of the linear alignment model are changes in shape: Both ellipticals and spirals should tend to be larger in low-density regions. Flexion-like distortions from higher-order tidal shear fields on galactic discs should exist but would be very small, given the result that the effect induced by $\partial^{3}\Phi$ on ellipticals as investigated in Giesel et al. (2022a) is already small. But physically, these distortions would be able to warp a galactic disc, which would constitute a link between alignments and disc dynamics.
@@
    Ellipticity spectra resulting from the linear alignment model naturally resemble those of elliptical galaxies as they are related just by a constant prefactor. With their smaller amplitudes, ellipticity correlations of spiral galaxies evade constraints on alignments derived from current weak lensing data and do not produce signals that can be measured at reasonable significances. More important, though, are the much smaller amplitudes of the linear alignment model in spiral galaxies compared to angular-momentum based models: Typically, the spectra are smaller by more than an order of magnitude at high multipoles. Additionally, the phenomenology of the spectra is quite different: In contrast to the quadratic, angular-momentum based alignment model which predicts no cross-correlation between the shapes of elliptical and spiral galaxies, the linear alignment model would predict cross-correlations between the shapes of the two galaxy types in a very natural way. Moreover, the linear alignment model would produce pure $E$-mode correlations in the ellipticity field, which would apply likewise for spiral and elliptical galaxies. Again, the quadratic model would be markedly different, with a high amplitude $B$-mode correlation being predicted. Lastly, one would expect a nonzero cross-correlation between intrinsic shapes of both ellipticals and spirals with gravitational lensing shear, with the typical anti-correlations present in both galaxy types. Linking tidal alignment to galaxy properties one should expect a dependence on Sérsic-index as more centrally concentrated objects would show weaker alignments. Estimates suggest that all intrinsic ellipticity spectra are within reach of Euclid. As outlined in Bretonnière et al. (2023), Sérsic profiles will play a crucial role in distinguishing galaxy morphologies in Euclid.
Spiral galaxies regularly have galactic bulges at their centres, which show analogous alignments in simulations by Jagvaral et al. (2022), up to the fact that they tend to have rather high Sérsic indices close to $n=4$, are rather strongly bound with high velocity dispersions $\sigma$ and are rather compact with a small Sérsic radii $R$; numbers of $\sigma=140~{}\mathrm{km}/\mathrm{s}$ and $R=1/2~{}\mathrm{kpc}$ have been reported for Milky Way (Valenti et al., 2018). Their amplitude seems to be similar to that of elliptical galaxies, and is much stronger than that of the host disc galaxy. Whether this is a consequence of the homologous distortion of virial systems remains to be investigated. We intend to develop the topic of tidally distorted galactic discs and the resulting ellipticity correlations further by considering the full 3-dimensional tidal field, which would effectively bridge between the concept of disc warps and intrinsic alignments, or perhaps even link alignments with tidally-induced bar formation in spiral galaxies.